solving systems by substitution
solve each system of equations using the substitution method. Check the solution.
x-7y=4
3x+y=-10

Answer by mananth(16946)

x-7y=4
3x+y=-10
x-7y=4
x=7y+4

3x+y=-10
3(7y+4)+y=-10
21y+12+y=-10
22y=-22
/22
y=-1
plug y in first equation
x-7y=4
x-7(-1)=4
x+7=4
x=-7+4
x=-3
x=-3
y=-1
Plug x& y in both equations to verify your answer